# The Quantum Measurement Problem: A Review of Recent Trends

### v1

@@ 1 Introduction
Quantum mechanics, as we learn in college, poses that an isolated system evolves unitarily and deterministically following the Schrödinger Equation. When we measure the system, the Born rule determines the output probabilities. The quantum state collapses in a non-unitary, stochastic way, so the same outcome is obtained if the measurement is repeated. Despite its awkward double-evolution standard, this seemingly simple set of rules enabled comprehension of nature and technological advances to a level our species has never seen.
Nevertheless, let’s ask what “measure the system” means. We will not find an answer in the standard quantum mechanics formulation. This innocent question unravels a disturbing thread of conceptual problems, which are surprisingly difficult to address experimentally and may have dramatic philosophical consequences over our conception of the world.
@@
The cacophony of proposed solutions for the Measurement Problem is disturbing. It hurts our conception of science as delivering a consistent description of reality. Why, after a century of quantum mechanics, do we still have no consensus on its most fundamental process? What do we truly understand about the Quantum Measurement Problem?
@@
Any newcomer trying to get an answer to these questions faces a formidable challenge: the Measurement Problem research is spread throughout uncountable isolated sub-communities. These groups don’t dialogue with each other, making it exceptionally difficult to build an overview of what is going on. The large number of single-author papers doesn’t help either. These papers raise hypotheses and propose theories and tests. However, most don’t resonate within a mature community and only add to the cacophony.
@@
As newcomers in the field ourselves—we just published our first contribution recently [1]—we felt firsthand all these issues. We tackled this challenge with extensive reading, discussion, and consult with specialists on specific topics. We explored all corners of the field rather than focusing on a single subfield. This review arose as a natural consequence. Our goal is to deliver a broad overview of the several competing research lines. We don’t aim at a comprehensive survey (which should not be possible within reasonable limits of a paper) but to illuminate the main questions being currently discussed, allowing curious non-specialists to start digging the literature by themselves. In doing so, we hope to show that the Measurement Problem is not as bleak as it seems, although there is still much to uncover.
@@
Given the Measurement Problem’s central role, it’s unsurprising that several other reviews have been recently published to explore its unresolved aspects from different perspectives. Hance and Hossenfelder [2] critically assess whether the problem is fundamental or an artifact of our theoretical framework. In a synthetic review, they identify five requirements that any satisfactory solution to the problem must satisfy and analyze several proposed solutions. Müller [3] followed a different path. Instead of focusing on the proposed solutions, he approaches the measurement problem from a formal logic perspective. He decomposes it into six independent problems, which should be investigated separately. Bassi, Dorato, and Ulbricht [4] provide an in-depth analysis of objective collapse models, discussing them from theoretical, experimental, and philosophical perspectives. It perfectly updates Bassi and co-workers’ excellent (and extensive) review published in 2013 [5]. Meanwhile, Freire Jr. [6] delivers a historical account, linking the foundations of quantum mechanics to current scientific and technological advances.
@@
Each of these works is worth the attention of readers interested in the status of foundational discussions of quantum mechanics. Nevertheless, none deliver a broad coverage of the Measurement Problem; that’s precisely what we missed when we started our journey. Addressing this gap, our review organizes the various proposed solutions into a structured overview, offering a broad reference point for both specialists and newcomers.
@@
Despite being unconventional in scientific papers, we extensively referred to non-peer-reviewed web resources, including books, essays, and videos. These resources offer plain language introduction and analysis without intimidating technicalities. They naturally don’t replace technical literature but should not be neglected as relevant entry points to the subject.
@@ 2 Quantum State Description and Evolution
Quantum mechanics encodes all information about a physical system in the quantum state , a complex-valued vector in the Hilbert Space containing contributions from all potential measurement outcomes.
@@
Consider, for example, the ammonia molecule NH3 (Figure 1). Its more stable geometry resembles a pyramid, with the three hydrogen atoms forming a triangular basis and the nitrogen atom lying above or below it. Thus, ignoring molecular vibrations and rotations, the nuclear quantum state of ammonia has two components, corresponding to pyramids up and down [13, Chs 8 & 9]. Let’s name them and states, respectively, which means that we will treat NH3 as a quantum bit (or a qubit).
@@
In this example, the general quantum state is a bi-dimensional vector (see Figure 2) written as
@@
where and are complex-valued coefficients (or amplitudes) with .
Eq.(1) trivially generalizes to any number of dimensions as . For a continuous basis, it becomes , where the amplitudes are called wave functions.
is usually a dynamic quantity evolving according to the Schrödinger Equation
where is the Hamiltonian operator. Alternatively, the dynamics of can be described in terms of the time evolution of the amplitudes, as indicated in Figure 2(c).
When we measure the system, we always observe a single outcome, an eigenvalue of the Hermitian operator representing the observable. The probability of each outcome is the square of the modulus of the corresponding state amplitude at the measurement time . That means,
When we determine the outcome value, we say the quantum state collapsed.
Alternatively, we can describe the collapse in terms of projector operators. In our example, the operators are
Suppose that the molecule is initially in state given by Eq. (2) at the measurement time . If we project on state , we have the trivial result
Thus, the probability of getting is
and the state after measurement is
Naturally, similar equations hold for measuring . Eqs.(6) and (7) are entirely equivalent to Eq.(3) (apart from a global phase) but only refer to the projection operator, not the amplitudes. The advantage of this formalism, called projection-valued measure, is that it describes the collapse that can be generalized to ensembles of states [14].
@@
All this discussion about collapse has assumed that it is instantaneous and occurs at tm. Standard quantum theory gives no clue on whether this is precise or not. Recent resonant inelastic X-ray scattering (RIXS) experiments have suggested that state collapse occurs gradually over a finite time [15], but this interpretation remains ambiguous, as the observed dynamics could also be explained by progressive decoherence rather than a fundamental modification of quantum measurement. We will return to when and how long collapse takes place in the discussion of Objective Collapse Theories in Section 5.3. The implications of an instantaneous collapse are explored in Section 5.7, in the context of quantum field measurements.
@@ 3 The Problems with Quantum Measurements
To summarize the description of the quantum state evolution discussed in the previous section, standard quantum mechanics prescribes that a quantum state follows two types of time evolution: Left on itself, it evolves with the Schrödinger Equation, which is unitary and deterministic [16, Ch 3][17, Ch 1]. During a measurement, however, the quantum state evolves with the Born rule, which is non-unitary and stochastic.
There are different ways to scrutinize this description [3, 18, 19]. Here, we follow Schlosshauer [20, Ch 2], who decomposed it into three problems: the problem of the preferred basis, the problem of the nonobservability of interference, and the problem of outcomes. Each of these challenges arises naturally from the standard formulation of quantum mechanics and highlights a gap in our understanding of what happens when a quantum system is measured. Below, we define these problems in sequence.
The problem of the preferred basis: Suppose we have an apparatus to detect whether the NH3 molecule is up or down. This apparatus is also supposed to be a quantum system, which is initially in the state “ready.” If it measures the molecule up, its pointer moves to 1. If it detects down, it moves to 0:
@@
the interaction with the apparatus will lead to the entangled state
Note that we assumed the interaction does not alter the system’s state. This one-to-one correlation between the system and the measurement apparatus establishes the von Neumann measurement scheme [20], which now only depends on quantum state collapse to determine the outcome.
We are free to change the basis representation in quantum mechanics. The states of the molecule can also be described, for example, on the following basis (Figure 3):
which has a one-to-one association with the apparatus basis:
In this new basis, the premeasured system-apparatus state is
Thus, it would appear that once the apparatus measured the molecule in the basis, the measurement in the (or any other arbitrary basis) is automatically determined. This contradicts the fundamental principle of quantum mechanics, as the observables related to and are non-commutative (note that the previous relations are equivalent to the Pauli matrices and ) and cannot be simultaneously well-defined. Moreover, this description is also unsettling, as the apparatus set to measure is physically distinct from a setup to measure .
We know that in practice, the apparatus designed to measure will produce outputs on this basis. Therefore, we must explain what eliminates basis ambiguity in quantum mechanics formalism, selecting a specific basis. This is the problem of the preferred basis, first spelled out by Zurek [21].
The problem of the nonobservability of interferences: Quantum mechanics predicts that systems can exist in coherent superpositions of states, and interference effects should be observable, such as interference patterns in double-slit experiments. However, in everyday macroscopic systems, we do not see such interference. Instead, objects appear to have well-defined classical properties. Why do we not see superpositions everywhere?
The ammonia molecule from our example may, in principle, exist as a superposition of its ’up’ and ’down’ states. Still, measurements always find it in one or the other, never in a visibly interfering combination. We may feel tempted to attribute it to tiny de Broglie wavelengths of large systems. However, this limitation can be circumvented, for instance, by using effects such as Talbot diffraction [22]. Under carefully controlled experimental conditions, interference patterns have been reported for molecules as large as 2000 atoms [23]. However, it requires isolation from environmental disturbances never observed in ordinary mesoscopic systems. This raises the fundamental question of how the environment suppresses interferences.
@@
The problem of outcomes: Even after addressing the previous two issues—identifying a preferred basis and suppressing interference—there remains a fundamental question: why do measurements yield definite outcomes rather than superpositions? The final state of the system-apparatus interaction is still a quantum superposition of different possible results, yet we never observe such superposition.
Furthermore, even after a basis is selected and interference is suppressed, quantum mechanics does not explain why a specific outcome occurs in each measurement. The Born rule successfully predicts the statistical distribution of outcomes over many repeated experiments, but it does not explain why a particular result is obtained in a single instance.
@@ 4 Decoherence
The problem of the nonobservability of interferences is addressed by decoherence. Decoherence disperses the quantum information about a system into correlations with the surrounding environment, suppressing the observable effects of superposition [20, 24]. To have a feeling of how decoherence acts, suppose that our and system can interact with an environment with only three states, , , and . We initially prepare the system in a superposition of and and the environment in , such that the initial state of the system plus environment is
@@
In general, the environmental states are not orthogonal, and the coherences are non-null. However, as time passes, the environment states tend to become orthogonal because their dynamics are different when the environment interacts with the system’s and geometries. Thus, the off-diagonal terms tend to zero
@@
In formal terms, decoherence arises from the time propagation of the reduced density matrix, which evolves with a master equation of the type [20, Ch 4].
where is the system’s Hamiltonian, including the environment’s perturbation, and is the term responsible for decoherence to the environment.
Decoherence tends to be extremely fast when considering the environment as a multidimensional system. For example, we can employ the Joos-Zeh model [25] to estimate the decoherence time of the and superposition in ammonia gas. According to this model, the coherence in Eq.(17) decays exponentially
@@
Although decoherence typically occurs in femtosecond or even sub-femtosecond timescales, specific molecular systems exhibit surprisingly long-lived coherence effects. Kaufman et al. [26], for instance, demonstrated that electronic coherences could survive for hundreds of femtoseconds in molecules with parallel potential energy surfaces due to reduced dephasing between branched wavepackets (see also 4). Schultz et al. [12] provide a comprehensive review of cases where vibronic coupling and structured environments lead to long coherence lifetimes, including examples from light-harvesting complexes in photosynthesis.
Babcock et al. [27] reported superradiant effects in tryptophan networks as large as microtubules, which require long-lived coherent effects. Their result is astonishing, given that fluorescence requires nanoseconds, and it isn’t easy to rationalize how coherence could survive for such a long time. For comparison, Tegmark estimated the decoherence time in microtubules to be shorter than 10-13 s [28]. Babcock’s result has also attracted media attention (see here and here) because it may support Hammeroff-Penrose’s hypothesis on the quantum origin of consciousness (see Section 5.6.2).
@@
In addition to suppressing quantum interferences, decoherence also addresses the problem of the preferred basis. It leads to the selection of a stable outcome basis through a process known as einselection, which drives the quantum state to the basis least entangled with the environment [21]. This preferred basis depends on the Hamiltonian [30]. Assuming that the position operator commutes with the system-environment interaction Hamiltonian (which is usually the case), decoherence tends to select the position eigenstates as the preferred basis as long as the interaction Hamiltonian is more important for the dynamics than the system’s Hamiltonian. If both terms compete, the preferred basis is close to coherent states, which minimizes the position and momentum uncertainty. If the system’s Hamiltonian dominates the dynamics, energy eigenstates are preferred as the basis.
For a molecule in a gas, Paz and Zurek analysis [30] implies that the outcome basis the environment selects for measuring nuclei is approximately position states. On the other hand, energy states emerge as the preferred outcome basis for measuring electrons. Thus, when we initially chose the basis for describing the quantum state of the ammonia molecule, we did it because it matched our classical intuition of nuclei corresponding to particles with defined positions. After discussing decoherence, we realize that we may have developed such a classic intuition because the environment selects the basis and suppresses superposition effects, creating the world of particles with well-defined positions.
Nevertheless, einselection may not be the last word in the preferred basis problem. Adil and -coauthors have recently shown that the preferred basis is not solely dictated by the system-environment interaction but also by the natural factorization of the global Hilbert space [31]. Different factorizations can lead to multiple coexisting sets of pointer states, meaning classical behavior emerges in different ways depending on the structure of the Hamiltonian. These co-existing classical realms challenge the idea that einselection always yields a single classical outcome. They may even reignite the preferred basis problem, but now at the level of Hamiltonian factorization.
@@
This dominant role of decoherence in the current understanding of quantum mechanical processes doesn’t imply it is a consensus. Landsmann, for instance, characterizes decoherence as “an unmitigated disaster” [32, Ch 11]. His core criticism is that decoherence rigorously only occurs in an environment with infinite degrees of freedom after infinite interaction time with the system. Kastner reinforces this criticism by noticing that decoherence claims of describing irreversible processes suffer from similar conceptual problems already present in statistical mechanics when explaining the emergence of irreversibility from reversible processes [33].

### v2

@@ 1 Introduction
Quantum mechanics, as we learn in college, posits that an isolated system evolves unitarily and deterministically following the Schrödinger Equation. When we measure the system, the Born rule determines the output probabilities. The quantum state collapses in a non-unitary, stochastic way, so the same outcome is obtained if the measurement is repeated. Despite its awkward double-evolution standard, this seemingly simple set of rules enabled comprehension of nature and technological advances to a level our species has never seen.
Nevertheless, let us ask what “measure the system” means. We will not find an answer in the standard quantum mechanics formulation. This innocent question unravels a disturbing thread of conceptual problems, which are surprisingly difficult to address experimentally and may have dramatic philosophical consequences for our conception of the world.
@@
The cacophony of proposed solutions to the Measurement Problem—including voices from parts of the physics community that question whether the problem even exists—is disturbing. It hurts our conception of science as delivering a consistent description of reality. Why, after a century of quantum mechanics, do we still have no consensus on its most fundamental process? What do we truly understand about the Quantum Measurement Problem?
@@
Any newcomer trying to get an answer to these questions faces a formidable challenge: the Measurement Problem research is spread throughout countless isolated sub-communities. These groups don’t dialogue with each other, making it exceptionally difficult to build an overview of what is going on. The large number of single-author papers doesn’t help either. These papers raise hypotheses and propose theories and tests. However, most don’t resonate within a mature community and only add to the cacophony.
@@
As newcomers in the field ourselves—we just published our first contribution recently [1]—we felt firsthand all these issues. We tackled this challenge with extensive reading, discussion, and consultation with specialists on specific topics. We explored all corners of the field rather than focusing on a single subfield. This review arose as a natural consequence. We aim to deliver a broad overview of the several competing research lines. We don’t aim at a comprehensive survey (which should not be possible within reasonable limits of a paper) but to illuminate the main questions being currently discussed, allowing curious non-specialists to start digging the literature by themselves. In doing so, we hope to show that the Measurement Problem is not as bleak as it seems, although there is still much to uncover.
@@
Given the Measurement Problem’s central role, it’s unsurprising that several other reviews have been recently published to explore its unresolved aspects from different perspectives. Hance and Hossenfelder [2] critically assess whether the problem is fundamental or an artifact of our theoretical framework. In a synthetic review, they identify five requirements that any satisfactory solution to the problem must satisfy and analyze several proposed solutions. Müller [3] followed a different path. Instead of focusing on the proposed solutions, he approaches the measurement problem from a formal logic perspective. He decomposes it into six independent problems, which should be investigated separately. Meanwhile, Freire Jr. [4] delivers a historical account, linking the foundations of quantum mechanics to current scientific and technological advances.
Surveys dedicated to objective collapse theories are numerous. Bassi, Dorato, and Ulbricht [5] present an in-depth analysis of objective collapse models from theoretical, experimental, and philosophical perspectives. Carlesso et al. [6] further explore experimental tests beyond interferometric setups, while Donadi and Bassi [7] review gravity-induced collapse models. These studies update the comprehensive review by Bassi and collaborators from 2013 [8] and build on the foundational report on dynamical reduction models by Ghirardi and Bassi [9], which remains a central reference in the field.
@@
Each of these works is worth the attention of readers interested in the status of foundational discussions of quantum mechanics. Nevertheless, none delivers a broad coverage of the Measurement Problem; that’s precisely what we missed when we started our journey. Addressing this gap, our review organizes the various proposed solutions into a structured overview, offering a broad reference point for both specialists and newcomers.
@@
Despite being unconventional in scientific papers, we extensively referred to non-peer-reviewed web resources, including books, essays, and videos. These resources offer a plain language introduction and analysis without intimidating technicalities. They naturally don’t replace technical literature, but should not be neglected as relevant entry points to the subject.
@@ 2 Quantum State Description and Evolution
This section briefly introduces basic quantum mechanics concepts, focusing on the elements most relevant to the following discussions. For in-depth treatments, we refer the reader to standard textbooks [16, 17]. Concise and accessible overviews can be found in Refs. [18, 19].
@@
Quantum mechanics encodes all information about a physical system in the quantum state , a complex-valued unit vector in the Hilbert Space . In a finite Hilbert space, all unit vectors are possible physical states of the system.
@@
Consider, for example, the ammonia molecule NH3 (Figure 1). Its more stable geometry resembles a pyramid, with the three hydrogen atoms forming a triangular basis and the nitrogen atom lying above or below it. Thus, ignoring molecular vibrations and rotations, the nuclear quantum state of ammonia has two components, corresponding to pyramids up and down [20, Chs 8 & 9]. Let’s name them and states, respectively, meaning we will treat NH3 as a quantum bit (or a qubit).
@@
In this example, the general quantum state is a two-dimensional vector (see Figure 2-a) written as
@@
where and are complex-valued coefficients (or amplitudes) with . Eq.(1) trivially generalizes to any number of dimensions as
For a continuous basis, it becomes , and the amplitudes are called wave functions.
So far, our example has been restricted to unit vectors, also known as pure states. Nonetheless, the usual situation is imperfect knowledge, where we only know that the system is in one of several possible pure states , each occurring with probability . In this case, it is more appropriate to describe the system using a density operator
which captures statistical uncertainty over an ensemble of pure states. Densities are operators with unit trace and satisfy for any normalized state , regardless of the chosen basis. A density operator represents a pure state if ; otherwise, the state is mixed.
For a composite system consisting of two subsystems and with Hilbert spaces and , the total Hilbert space is the tensor product
A pure state of the total system is a unit vector in . It is called separable if it can be written as
otherwise, it is said to be entangled. and are normalized, not necessarily orthogonal, states of the two systems. A general pure state can be expressed as a superposition of product basis states:
For mixed states, a densityoperator is said to be separable if it can be written as
where each and is a density operators on and , and , . Otherwise, is entangled.
The generalization to more than two subsystems is straightforward.
The state vector is a dynamical quantity that evolves unitarily according to , which implies it satisfies the Schrödinger equation:
where is the Hamiltonian operator (see Figure 2-b). Equivalently, the density operator evolves according to the von Neumann equation:
where denotes the commutator.
Every measurable physical quantity (observable) is represented by a self-adjoint operator on . The possible outcomes of a measurement of are its eigenvalues . For a non-degenerate spectrum, the probability of obtaining the outcome at measurement time is (see Figure 2-c)
In terms of the density operator, this probability becomes
This formulation corresponds to a projective measurement (also known as a projection-valued measure, or PVM), where the outcome is associated with the projection operator . For more general types of quantum measurements—such as those involving degenerate spectra or positive operator-valued measures (POVMs)—we refer the reader to Ref. [19] for a pedagogical overview.
@@
According to the standard postulates of quantum mechanics, after the measurement, the system is left in the normalized eigenstate if was measured. This state change is referred to as quantum state collapse. In this discussion, we assumed that collapse is instantaneous and occurs at . Standard quantum theory gives no clue on whether this is precise or not. Recent resonant inelastic X-ray scattering (RIXS) experiments have suggested that state collapse occurs gradually over a finite time [21]. However, this interpretation remains ambiguous, as the observed dynamics could also be explained by progressive decoherence rather than a fundamental modification of quantum measurement. We will return to when and how long it takes for collapse to occur in discussing Objective Collapse Theories in Section 5.3. The implications of an instantaneous collapse are explored in Section 5.7, in the context of quantum field measurements.
@@ 3 The Problems with Quantum Measurements
To summarize the description of the quantum state evolution discussed in the previous section, standard quantum mechanics prescribes that a quantum state follows two types of time evolution: Left on itself, it evolves with the Schrödinger Equation, which is unitary and deterministic [16, Ch 3][17, Ch 1]. During a measurement, the system produces a classical outcome with probabilities given by the Born rule. In idealized cases, this is often described by a non-unitary state update (the projection postulate), but more generally, a measurement is any transformation that extracts classical information from a quantum system. These processes must be probabilistic and cannot allow communication between distant systems without a physical carrier.
There are different ways to scrutinize this description [3, 22, 23]. Here, we follow Schlosshauer [24, Ch 2], who decomposed it into three problems: the problem of the preferred basis, the problem of the nonobservability of interference, and the problem of outcomes. Each of these challenges arises naturally from the standard formulation of quantum mechanics and highlights a gap in our understanding of what happens when a quantum system is measured. Below, we define these problems in sequence.
The problem of the preferred basis. Suppose we have an apparatus to detect whether the NH3 molecule is up or down. This apparatus is also modeled as a quantum system, initially in a “ready” state. If it measures the molecule up, its pointer moves to 1; if it detects down, it moves to 0 (in the following, we will omit for brevity):
@@
the measurement interaction leads to an entangled state:
This one-to-one correlation between system and apparatus defines the premeasurement state in the von Neumann measurement scheme [24], where a collapse (or equivalent mechanism) is needed to produce a definite outcome.
The same final state can be expressed in a different basis. Define (Figure 3):
and
Then the same entangled state becomes
Formally, both descriptions are equivalent. But in practice, the apparatus always produces outcomes on a specific basis—here, —not in or any other. One might argue that the apparatus, by design, selects the measurement basis. But this assumes what needs to be explained: how and why a particular basis emerges from the unitary quantum dynamics. After all, the entangled state can be expressed in many incompatible bases, and the Schrödinger equation itself does not single out any of them.
This ambiguity, why a specific set of outcomes is realized when many are in principle available, is the problem of the preferred basis, first formulated by Zurek [25].
The problem of the nonobservability of interference. Quantum mechanics predicts that systems can exist in coherent superpositions, which should produce observable interference effects. Indeed, such effects are routinely seen at microscopic scales—for instance, in electron double-slit experiments [26] or the oscillations of generalized oscillator strengths for symmetric molecules under electron impact [27]. More recently, Kanitz and co-workers demonstrated interference patterns from helium and hydrogen atoms transmitted through single-layer graphene, achieving diffraction with atoms at kiloelectronvolt energies [28].
Nevertheless, interference is never observed in macroscopic objects: tables, measurement devices, or Schrödinger’s cat do not display interference fringes. One might suspect this absence is simply due to the extremely short de Broglie wavelengths of large systems, making interference effects undetectable. However, this explanation is insufficient. Under carefully controlled conditions, interference remains observable even for massive molecules with minuscule de Broglie wavelengths, as demonstrated by Talbot diffraction experiments with molecules comprising up to 2000 atoms [29]. Thus, the absence of interference in everyday situations cannot be entirely attributed to technical limitations.
The real puzzle is why coherence, though allowed in principle by the Schrödinger equation, becomes unobservable in practice beyond microscopic scales. This disconnect defines the problem of the nonobservability of interference and points toward the need for a dynamical mechanism that suppresses coherence in macroscopic systems.
@@
The problem of outcomes. Even after addressing the previous two issues—identifying a preferred basis and suppressing interference—a fundamental question remains: Why do measurements yield a single, definite outcome rather than a superposition of possible results? After decoherence, the total system (including apparatus and environment) is still described by an entangled quantum state encompassing all possible outcomes (as shown in Eq.14. Yet we never observe such superpositions: each measurement yields a unique result. Indeed, what would it even mean to observe a superposition?
Standard quantum mechanics provides no mechanism for this selection. The Born rule accurately predicts the statistical distribution of outcomes over many measurements. Still, it does not explain why a particular result occurs or what selects it in an individual run.
@@ 4 Decoherence
The problem of the nonobservability of interference is addressed by decoherence. Decoherence disperses the quantum information about a system into correlations with the surrounding environment, suppressing the observable effects of superposition [24, 30]. To have a feeling of how decoherence acts, suppose that our and system can interact with an environment with only three states, , , and . We initially prepare the system in a superposition of and and the environment in , such that the initial state of the system plus environment is
@@
In general, the environmental states are not orthogonal, and the coherences are non-null. However, as time passes, the environment states tend to become orthogonal because their dynamics are different when the environment interacts with the system’s and geometries. Thus, the off-diagonal terms tend to zero,
@@
In formal terms, decoherence arises from the time propagation of the reduced density matrix, which evolves with a master equation of the type [24, Ch 4].
where is the system’s Hamiltonian, including the environment’s perturbation, and it is the term responsible for decoherence to the environment.
Decoherence tends to be extremely fast when considering the environment as a multidimensional system. For example, we can employ the Joos-Zeh model [31] to estimate the decoherence time of the and superposition in ammonia gas. According to this model, the coherence in Eq.(21) decays exponentially,
@@
Although decoherence typically occurs in femtoseconds or even sub-femtoseconds, specific molecular systems exhibit surprisingly long-lived coherence effects. For instance, Kaufman et al. [32] demonstrated that electronic coherences could survive for hundreds of femtoseconds in molecules with parallel potential energy surfaces due to reduced dephasing between branched wavepackets (see also 4). Schultz et al. [15] comprehensively review cases where vibronic coupling and structured environments lead to long coherence lifetimes, including examples from light-harvesting complexes in photosynthesis.
Babcock et al. [33] reported superradiant effects in tryptophan networks extending over microtubule-scale architectures, which require long-lived electronic coherence. This is remarkable, since superradiance typically unfolds over hundreds of femtoseconds [34]. For isolated tryptophan in aqueous solution, electronic decoherence occurs in less than 100 fs [35], implying that the protein scaffold must be critical in preserving coherence to allow superradiant emission. Babcock’s result has also attracted media attention (see here and here) because it may support Hammeroff-Penrose’s hypothesis on the quantum origin of consciousness (see Section 5.6.2).
@@
In addition to suppressing quantum interferences, decoherence also addresses the problem of the preferred basis. It leads to the selection of a stable outcome basis through a process known as einselection, which drives the quantum state to the basis least entangled with the environment (this basis is known as the pointer basis) [25]. This preferred basis depends on the Hamiltonian [37]. Assuming that the position operator commutes with the system-environment interaction Hamiltonian (usually the case), decoherence tends to select the position eigenstates as the preferred basis as long as the interaction Hamiltonian is more important for the dynamics than the system’s Hamiltonian. If both terms compete, the preferred basis is close to coherent states, which minimizes the position and momentum uncertainty. If the system’s Hamiltonian dominates the dynamics, energy eigenstates are preferred as the basis.
For a molecule in a gas, Paz and Zurek’s analysis [37] implies that the outcome basis of the environment selects for measuring nuclei is approximately position states. On the other hand, energy states emerge as the preferred outcome basis for measuring electrons. Thus, when we initially chose the basis for describing the quantum state of the ammonia molecule, we did it because it matched our classical intuition of nuclei corresponding to particles with defined positions. After discussing decoherence, we realize that we may have developed such a classic intuition because the environment selects the basis and suppresses superposition effects, creating the world of particles with well-defined positions.
Nevertheless, einselection may not be the last word in the preferred basis problem. Adil and co-authors have recently shown that the preferred basis is not solely dictated by the system-environment interaction but also by the natural factorization of the global Hilbert space [38]. Different factorizations can lead to multiple coexisting sets of pointer states, meaning classical behavior emerges in different ways depending on the structure of the Hamiltonian. These coexisting classical realms challenge the idea that einselection always yields a single classical outcome. They may even reignite the preferred basis problem, but now at the level of Hamiltonian factorization.
@@
This dominant role of decoherence in the current understanding of quantum mechanical processes doesn’t imply it is a consensus. Landsmann, for instance, characterizes decoherence as “an unmitigated disaster” [39, Ch 11]. His core criticism is that decoherence rigorously only occurs in an environment with infinite degrees of freedom after infinite interaction time with the system. Kastner reinforces this criticism by noticing that decoherence claims of describing irreversible processes suffer from similar conceptual problems already present in statistical mechanics when explaining the emergence of irreversibility from reversible processes [40].